# Sparse-DeRF: Deblurred Neural Radiance Fields from Sparse View

### v1

@@ V-F5 Inconsistency of Pre-deblurred Images
As mentioned in Section I and IV-C, image deblurring is conducted independently for each image and presents inconsistent geometry in various regions due to its ill-posed property. In Figure 11, we present the qualitative comparison to reveal the inconsistency of pre-deblurred images across the multi-view training images. Pre-deblurred images are acquired by applying the MPRNet [56]. In addition, we attach the reference images, which are rendered from the DP-NeRF [2] trained with full view, to help readers better understand the inconsistency issue and compare the pre-deblurred geometry to approximated ground truth geometry. Comparing emphasized regions from each image, the pre-deblurred image shows relatively well-restored textures in each image, but the geometry is inconsistently restored and distorted across the multiple views. Such inconsistency adversely affects the learning of radiance fields since it is trained by pixel-wise color reconstruction loss. Furthermore, the performance of image deblurring varies even within a single image depending on the structural complexity of the local region, making it challenging to learn blur kernels. Due to these reasons, directly utilizing pre-filtered images for training radiance fields is difficult. Therefore, perceptual distillation is introduced to transfer only the perceptual texture of pre-deblurred images to the radiance fields.
@@ VI Limitations and Discussions
Despite the remarkable enhancement in terms of 3D geometry and appearance, there are still several limitations. The first one is derived from the blur kernel itself, especially the relationship between the type of the blur kernel and the properties of each scene. For example, the performance is more improved with the rigid blur kernel of DP-NeRF [2] in some scenes, but in other scenes, the improvements are greater with the flexible blur kernel of Deblur-NeRF [1]. These kernel-dependent performances are different across the scenes. As we figure out, a flexible kernel leads to reduced space ambiguity but high scene distortion. In contrast, the rigid kernel leads to accurate geometry but suffers difficulty in optimizing the scene where the distances of the objects from the camera in the scene are diverse and some objects are located very close to the camera due to the inherent rigidity. We tried to take advantage of both kernels and design the hybrid kernel to maximize the effectiveness of the Sparse-DeRF, but it didn’t work as we imagined. Constructing the hybrid blur kernel that has rigid and flexible properties can be a promising future research direction regardless of sparse view setting in the deblurred neural radiance fields (DeRF). The second one is that we have to set the proper hyper-parameter for MGS to find the most effective function shape although MGS greatly improves the 3D geometry in the DeRF from sparse view. However, it is difficult to find an ideal function shape according to the arrangement of the object in the scene, especially as we mentioned above. We handle these cases by setting the magnitude $\rho$ as a high value to only ignore the gradient in a very near distance region, which is attached to the appendix as detailed hyper-parameters per each scene. In this sense, this manual setting of hyperparameters is regarded as one of our limitations. We believe the limitation can be alleviated in future research through various methods such as the introduction of learnable parameters for gradient scaling function. Finally, although sparse view setting of blurry inputs is an extremely practical scenario for blurry inputs, it is too hard to enhance the performance of the DeRF in the 2-view setting due to the lack of the scene information included in the input data. The innate challenge of the 2-view setting is that sparse overlapped 3D space usually leads to inaccurate geometry, which is more likely to be mapped to be painted texture on the wall at the near or far depth regions. There is still room to improve the visual quality of the Sparse-DeRF and solve these problems through state-of-the-art generative methods such as diffusion models, which can be a great future direction for constructing the DeRF from sparse view.
@@ Appendix A Sample Space Visualization
TABLE V: Image indices of each scene for training Sparse-DeRF.
@@
TABLE VI: Hyper parameters of modulated gradient scaling (MGS) for entire scenes of the Deblur-NeRF [1] dataset. $\rho$ and $\eta$ denote the magnitude and period of sine function, respectively.
@@
Appendix B Additional Implementation Details
B-A Training Scene Indices
For fair comparison in future research, we present the image indices of each scene for training the Sparse-DeRF in TABLE V. The indices are manually selected from the training images of each scene as we mentioned in the main paper. Note that, the indices of the 2-view and 4-view settings are subsets of the 6-view setting.
B-B Parameters for Entire Scenes
In TABLE VI, we present the hyper-parameters of MGS for entire scenes, which consist of period $\eta$ and magnitude $\rho$ of the sine function. As we indicate in Section IV-B in the main paper, we set the magnitude $\rho$ as high value to only ignore the gradient in very near distance regions for scenes such as Buick, Puppet, Cozyroom, Factory, Pool, and Trolley. Please refer to the figure of the function shape depending on the parameters in the main paper.
Appendix C Additional Quantitative Results
C-A Quantitative Results for Entire Scenes
We present the comprehensive experimental results for entire scenes from 2-view, 4-view, and 6-view settings in TABLE VIIVIII, and IX, respectively. The results of the Sparse-DeRF in the TABLE VIIVIII, and IX are representative results as we presented in Section V-E1 in the main paper. In addition, we present more detailed ablation results for entire scenes from 2-view, 4-view, and 6-view settings in TABLE XXI, and XII, respectively. As we mentioned before, we also present the separated ablation results according to the type of the kernel from DP-NeRF [2] and Deblur-NeRF [1].
C-B Complex Optimization Problem for Entire Scenes
We present the comprehensive experimental results of the complex optimization problem for entire scenes from 2-view, 4-view, and 6-view settings in TABLE XIVXV, and XVI, respectively. In addition, we also present the quantitative results of the experiments on the Decoration scene that varies the number of sparse views in TABLE XIII as we mentioned in the main paper.
C-C Comparison to Naive Gradient Scaling for Entire Scenes
We present the quantitative comparison results for entire scenes that compare our modulated gradient scaling (MGS) and naive gradient scaling of [51] from 2-view, 4-view, and 6-view settings in TABLEXVIIXVIII, and XIX, respectively.
TABLE VII: Quantitative results of novel view synthesis for the entire scene of synthetic and real scenes obtained from 2-view settings. Each color shading represents the best, second best and third best result, respectively. DP-kernel and DN-kernel denote the kernel of [2] and [1].
@@ C-C Comparison to Naive Gradient Scaling for Entire Scenes
TABLE VIII: Quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 4-view settings. Each color shading represents the best, second best and third best result, respectively. DP-kernel and DN-kernel denote the kernel of [2] and [1].
@@
TABLE IX: Quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 6-view settings. Each color shading represents the best, second best and third best result, respectively. DP-kernel and DN-kernel denote the kernel of [2] and [1].
@@
TABLE X: Ablation quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 2-view settings. Each color shading represents the best, second best and third best result, respectively. DP-kernel and DN-kernel denote the kernel of [2] and [1].
@@
TABLE XI: Ablation quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 4-view settings. Each color shading represents the best, second best and third best result, respectively. DP-kernel and DN-kernel denote the kernel of [2] and [1].
@@
TABLE XII: Ablation quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 6-view settings. Each color shading represents the best, second best and third best result, respectively. DP-kernel and DN-kernel denote the kernel of [2] and [1].
@@
TABLE XIII: Quantitative results of the RegNeRF [6] with and without the blur kernel from 2-view to 10-view settings. For comparison, the blur kernel of the DP-NeRF [2] is employed, which is denoted as DP-kernel. Color shading represents the better result.
@@
TABLE XIV: Entire experimental results for the complex optimization issue in both synthetic and real scenes from 2-view settings. Each color shading represents the best and second results for each experimental setting, respectively.
@@
TABLE XV: Entire experimental results for the complex optimization issue in both synthetic and real scenes from 4-view settings. Each color shading represents the best and second results for each experimental setting, respectively.
@@
TABLE XVI: Entire experimental results for the complex optimization issue in both synthetic and real scenes from 6-view settings. Each color shading represents the best and second results for each experimental setting, respectively.
@@
TABLE XVII: Quantitative comparison between naive gradient scaling of [51] and our MGS for entire scenes of synthetic and real scenes obtained from 2-view settings. The experimental results are presented according to the two types of kernel we utilize in this paper, which are kernels of the DP-NeRF [2] and Deblur-NeRF [1]. Each kernel is denoted as DP-kernel and DN-kernel, respectively. Color shading represents the better result.
@@
TABLE XVIII: Quantitative comparison between naive gradient scaling of [51] and our MGS for entire scenes of synthetic and real scenes obtained from 4-view settings. The experimental results are presented according to the two types of kernel we utilize in this paper, which are the kernels of DP-NeRF [2] and Deblur-NeRF [1]. Each kernel is denoted as DP-kernel and DN-kernel, respectively. Color shading represents the better result.
@@
TABLE XIX: Quantitative comparison between naive gradient scaling of [51] and our MGS for entire scenes of synthetic and real scenes obtained from 6-view settings. The experimental results are presented according to the two types of kernel we utilize in this paper, which are the kernels of DP-NeRF [2] and Deblur-NeRF [1]. Each kernel is denoted as DP-kernel and DN-kernel, respectively. Color shading represents the better result.

### v2

@@ V-F6 Inconsistency of Pre-deblurred Images
Image deblurring is conducted independently for each image and presents inconsistent geometry in various regions due to its ill-posed property. In Figure 13, we present the qualitative comparison to reveal the inconsistency of pre-deblurred images across the multi-view training images. Pre-deblurred images are acquired by applying the MPRNet [65]. In addition, we attach the reference images, which are rendered from the DP-NeRF [2] trained with full view, to help readers better understand the inconsistency issue and compare the pre-deblurred geometry to approximated ground truth geometry. Comparing emphasized regions from each image, the pre-deblurred image shows relatively well-restored textures in each image, but the geometry is inconsistently restored and distorted across the multiple views. Such inconsistency adversely affects the learning of radiance fields, which is crucial for perceptual quality, since it is trained by pixel-wise color reconstruction loss. In TABLE V, we present the comparison results of the NeRF [19], RegNeRF [6], and SparseNeRF [52] with and without the pre-deblurred images. The results from 4-view and 6-view also demonstrate that, it is not enough to reconstruct the clean radiance fields with high perceptual quality only depending on the pre-deblurring and depth estimation although they provide the abundant additional priors. To qualitatively demonstrate the observation, we included Figure 12, which compares the results trained on 4 view setting. While SparseNeRF [52] + MPR [65], which utilizes both pre-deblurred images and depth prior, achieved higher PSNR and SSIM scores numerically, the rendered results clearly show that our approach produces much sharper and more precise renderings. Due to these reasons, directly utilizing pre-deblurred images for training radiance fields is difficult. Therefore, perceptual distillation is introduced to transfer only the perceptual texture of pre-deblurred images to the radiance fields.
@@ VI Limitations and Discussions
Despite the remarkable enhancement in terms of 3D geometry and appearance, there are several points for improvement.
The first one is derived from the blur kernel itself, especially the relationship between the type of the blur kernel and the properties of each scene. For example, the performance is more improved with the rigid blur kernel of DP-NeRF [2] in some scenes, but in other scenes, the improvements are greater with the flexible blur kernel of Deblur-NeRF [1]. These kernel-dependent performances are different across the scenes. As we figure out, a flexible kernel leads to reduced space ambiguity but high scene distortion. In contrast, the rigid kernel leads to accurate geometry but suffers difficulty in optimizing the scene where the distances of the objects from the camera in the scene are diverse and some objects are located very close to the camera due to the inherent rigidity. We tried to take advantage of both kernels and design the hybrid kernel to maximize the effectiveness of the Sparse-DeRF, but it didn’t work as we imagined. Constructing the hybrid blur kernel that has rigid and flexible properties can be a promising future research direction regardless of sparse view setting in the deblurred neural radiance fields (DeRF).
The second one is that we have to set the proper hyper-parameter for MGS to find the most effective function shape although MGS greatly improves the 3D geometry in the DeRF from sparse view. However, it is difficult to find an ideal function shape according to the arrangement of the object in the scene. We handle these cases by setting the magnitude $\rho$ as a high value to only ignore the gradient in a very near distance region, which is attached to the appendix as detailed hyper-parameters per each scene. In this sense, this manual setting of hyperparameters is regarded as one of our limitations. We believe the limitation can be alleviated in future research through various methods such as the introduction of learnable parameters for gradient scaling function.
Finally, although sparse view setting of blurry inputs is an extremely practical scenario for blurry inputs, it is too hard to enhance the performance of the DeRF in the 2-view setting due to the lack of the scene information included in the input data. The innate challenge of the 2-view setting is that sparse overlapped 3D space usually leads to inaccurate geometry, which is more likely to be mapped to be painted texture on the wall at the near or far depth regions. There is still room to improve the visual quality of the Sparse-DeRF and solve these problems through state-of-the-art generative methods similar to the work like Diffusionerf [69], which can be a great future direction for constructing the DeRF from sparse view. On the other hand, designing models that account for inconsistency to utilize additional data-deriven priors, such as depth and normal priors from pretrained models, could be an interesting direction for future work. Furthermore, leveraging real-time rendering models, like the recently emerging 3D Gaussian Splatting [42], could enhance practical applicability.
@@ Appendix A Sample Space Visualization
TABLE VI: Average results of the comparison with the model that uses depth-prior. The experiments are conducted on both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings. Each color shading represents the best, second best, and third best results, respectively.
[ 2-view ] [ 4-view ] [ 6-view ]
SparseNeRF [52] 18.12 .4729 .4884 18.01 .4683 .5245 21.07 .6105 .4042 20.69 .5909 .4450 21.77 .6369 .3903 21.28 .6189 .4250
SparseNeRF [52] + Depth-Any V2 [68] 18.24 .4909 .4479 16.35 .4262 .4768 21.78 .6679 .3354 18.87 .5406 .4016 22.56 .7034 .3141 19.44 .5659 .3845
TABLE VII: Quantitative results for the entire scenes of real scenes from 2-view settings. Each color shading represents the best, second best, and third best result, respectively. Depth-Any V2 denotes the Depth-Anything V2 [68].
Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
SparseNeRF [52] 15.80 .3622 .5988 22.06 .7001 .3419 23.90 .6063 .3824 14.53 .3454 .5651 14.32 .3499 .5651 18.12 .4728 .4883
SparseNeRF [52] + Depth-Any V2 [68] 15.93 .3923 .5060 23.02 .7419 .2972 23.64 .5968 .3957 13.94 .3086 .5469 14.71 .4150 .4938 18.24 .4909 .4479
Real Scene Ball Basket Buick Coffee Decoration
SparseNeRF [52] 21.62 .5588 .5092 16.32 .4636 .5488 18.32 .5356 .4754 24.82 .8225 .2958 14.85 .3325 .6020
SparseNeRF [52] + Depth-Any V2 [68] 21.58 .5573 .5107 16.44 .4698 .5462 18.28 .5352 .4752 24.65 .8195 .3003 15.19 .3432 .5941
Real Scene Girl Heron Parterre Puppet Stair Average
SparseNeRF [52] 14.16 .4380 .5641 16.81 .3257 .5566 17.58 .3601 .5928 16.89 .4677 .5261 18.75 .3784 .5738 18.01 .4683 .5245
SparseNeRF [52] + Depth-Any V2 [68] 14.65 .4600 .5513 16.90 .3258 .5597 17.25 .3481 .5983 16.66 .4684 .5234 18.33 .3614 .5861 16.35 .4262 .4768
TABLE VIII: Quantitative results for the entire scenes of real scenes from 4-view settings. Each color shading represents the best, second best, and third best result, respectively. Depth-Any V2 denotes the Depth-Anything V2 [68].
Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
SparseNeRF [52] 17.15 .4271 .5516 23.50 .7531 .2994 26.63 .7091 .3114 18.93 .5838 .4389 19.14 .5865 .4195 21.07 .6105 .4042
SparseNeRF [52] + Depth-Any V2 [68] 18.93 .5813 .3959 24.86 .8044 .2335 26.74 .7047 .2988 18.90 .6111 .3955 19.46 .6379 .3482 21.78 .6679 .3354
Real Scene Ball Basket Buick Coffee Decoration
SparseNeRF [52] 21.56 .5539 .5020 20.45 .5986 .4620 21.04 .6392 .4008 26.55 .8499 .2739 18.94 .5128 .5042
SparseNeRF [52] + Depth-Any V2 [68] 21.74 .5587 .4973 20.55 .6007 .4611 21.17 .6474 .3987 26.47 .8504 .2696 18.86 .5097 .5031
Real Scene Girl Heron Parterre Puppet Stair Average
SparseNeRF [52] 17.96 .6579 .4445 19.08 .4873 .4793 20.91 .5239 .4808 18.56 .5409 .4444 21.83 .5446 .4574 20.69 .5909 .4450
SparseNeRF [52] + Depth-Any V2 [68] 18.10 .6633 .4371 19.03 .4856 .4807 21.07 .5276 .4777 18.67 .5506 .4378 21.99 .5530 .4545 18.87 .5406 .4016
TABLE IX: Quantitative results for the entire scenes of real scenes from 6-view settings. Each color shading represents the best, second best, and third best result, respectively. Depth-Any V2 denotes the Depth-Anything V2 [68].
Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
SparseNeRF [52] 17.17 .4286 .5402 23.86 .7601 .3021 28.04 .7612 .2772 19.99 .6216 .4212 19.78 .6131 .4106 21.77 .6369 .3903
SparseNeRF [52] + Depth-Any V2 [68] 19.42 .6130 .3777 25.10 .8130 .2282 28.09 .7628 .2643 19.78 .6529 .3686 20.45 .6751 .3318 22.56 .7034 .3141
Real Scene Ball Basket Buick Coffee Decoration
SparseNeRF [52] 22.17 .5654 .4939 21.86 .6733 .4081 21.63 .6719 .3747 26.17 .8382 .2874 19.94 .5696 .4646
SparseNeRF [52] + Depth-Any V2 [68] 22.14 .5661 .4896 21.96 .6746 .4099 21.66 .6754 .3736 26.21 .8373 .2847 20.04 .5763 .4615
Real Scene Girl Heron Parterre Puppet Stair Average
SparseNeRF [52] 18.92 .7073 .4091 19.09 .4705 .4895 21.54 .5630 .4536 19.65 .5951 .4017 21.79 .5345 .4673 21.28 .6189 .4250
SparseNeRF [52] + Depth-Any V2 [68] 19.40 .7156 .4050 19.17 .4731 .4900 21.68 .5701 .4500 19.80 .5995 .4000 21.84 .5369 .4657 19.44 .5659 .3845
Figure 15: Additional qualitative ablation results on Puppet, Ball, and Basket scene from 2-view, 4-view, and 6-view settings, which are the same scene in the main paper. The results show the comparison between SparseNeRF [52] and SparseNeRF [52] trained with the depth maps predicted from Depth-Anything V2 [68], which are indicated as reference number [52] and Depth-Any V2, respectively.
Appendix B Additional Analysis and Experimental Results
B-A Models with Depth Prior
We present the quantitative results of our model and SparseNeRF [52] in TABLE VIVIIVIII, and IX. Additionally, we include the results of SparseNeRF [52] trained with the more recent state-of-the-art monocular depth estimation model, Depth-Anything V2 [68], in these tables and present the qualitative comparison in Figure 15. Interestingly, the results show that SparseNeRF [52] using depth maps predicted by Depth-Anything V2 [68] yields lower PNSR and SSIM scores but a higher LPIPS score in real-world scenes compared to the originally adopted depth maps from DPT [70]. We believe that this degradation occurs because Depth-Anything V2 [68] predicts more accurate depth maps even from blurred images in real-world data. However, this improvement in depth estimation unexpectedly exacerbates multi-view inconsistency among the predicted depth maps, which in turn negatively affects the optimization process of the radiance field in real scenes. As a consequence, the multi-view inconsistency issue, which becomes particularly noticeable across four or more views, remains unresolved. This degradation is also evident in the lower quality of the rendered images in Figure 15.
TABLE X: Average results of MGS experiments on world coordinate systems. The experiments are conducted on both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings. Color shading represents the best and second results, respectively.
[ 2-view ] [ 4-view ] [ 6-view ]
DP-kernel $\times$ 14.46 .2614 .6258 13.26 .1797 .6732 18.46 .4554 .4691 16.92 .3679 .4570 19.04 .4822 .4408 19.38 .4760 .3740
DP-kernel + Naive [60] 14.41 .2579 .6204 13.20 .1789 .6680 18.40 .4506 .4706 17.05 .3654 .4544 19.01 .4830 .4425 19.12 .4575 .3831
DP-kernel + MGS 14.49 .2564 .5974 15.22 .2880 .5362 18.54 .4493 .4300 17.91 .4061 .4019 19.10 .4781 .3963 19.79 .4923 .3470
Figure 16: Additional qualitative ablation results on Basket, Coffee, and Decoration scene from 2-view, 4-view, and 6-view settings, which are trained on world coordinate system. The results demonstrate that our proposed MGS guides the geometry of the scenes and works properly even in the world coordinate system.
B-B Geometric Guidance of Modulated Gradient Scaling
We initially explored a gradient scaling function that could operate in normalized device coordinate (NDC). However, inspired by the idea that gradient scaling could serve as geometric guidance for the scene, we proposed MGS. To demonstrate that MGS can function as a structural clue regardless of the coordinate systems, we present TABLE X along with Figure 16. The experiments were conducted in the world coordinate system, using the same MGS parameters for each scene as presented in TABLE XI to match the scaling function shape in NDC for fair comparison. We converted the distances of ray samples along the rays in the world coordinate system to the NDC depth scale and used them as $\delta_{\textbf{s}_{i}}$ in Eq.21 of the main paper. Note that, other than this scaling adjustment, all settings for training the radiance fields remained in the world coordinate system. The kernel employed for the experiments is DP-kernel. The results show that our proposed MGS achieves superior performance even in the world coordinate system. These findings validate that MGS is not limited to a specific coordinate system but effectively captures the arrangement of scene components, facilitating the learning of radiance fields as intended. We additionally attach the quantitative results for entire scenes in Section D-D.
@@ B-B Geometric Guidance of Modulated Gradient Scaling
TABLE XI: Image indices of each scene for training Sparse-DeRF.
@@
TABLE XII: Hyper parameters of modulated gradient scaling (MGS) for entire scenes of the Deblur-NeRF [1] dataset. $\rho$ and $\eta$ denote the magnitude and period of sine function, respectively.
@@
Appendix C Additional Implementation Details
C-A Training Scene Indices
For fair comparison in future research, we present the image indices of each scene for training the Sparse-DeRF in TABLE XI. The indices are manually selected from the training images of each scene as we mentioned in the main paper. Note that, the indices of the 2-view and 4-view settings are subsets of the 6-view setting.
C-B Parameters for Entire Scenes
In TABLE XII, we present the hyper-parameters of Modulated Gradient Scaling (MGS) for entire scenes, which consist of period $\eta$ and magnitude $\rho$ of the sine function. As we indicate in Section IV-B in the main paper, we set the magnitude $\rho$ as high value to only ignore the gradient in very near distance regions for scenes such as Buick, Puppet, Cozyroom, Factory, Pool, and Trolley. Please refer to the figure of the function shape depending on the parameters in the main paper.
Appendix D Additional Quantitative Results
D-A Quantitative Results for Entire Scenes
We present the comprehensive experimental results for entire scenes from 2-view, 4-view, and 6-view settings in TABLE XIIIXIV, and XV, respectively. In the tables, DP-kernel and DN-kernel denote the kernel of [2] and [1], respectively. The results of the Sparse-DeRF in the TABLE XIIIXIV, and XV are representative results as we presented in Section V-E1 in the main paper. In addition, we present more detailed ablation results for entire scenes from 2-view, 4-view, and 6-view settings in TABLE XVIXVII, and XVIII, respectively. As we mentioned before, we also present the separated ablation results according to the type of the kernel from DP-NeRF [2] and Deblur-NeRF [1].
D-B Complex Optimization Problem for Entire Scenes
We present the comprehensive experimental results of the complex optimization problem for entire scenes from 2-view, 4-view, and 6-view settings in TABLE XXXXI, and XXII, respectively. In addition, we also present the quantitative results of the experiments on the Decoration scene that varies the number of sparse views in TABLE XIX as we mentioned in the main paper.
D-C Comparison to Naive Gradient Scaling for Entire Scenes
We present the quantitative comparison results for entire scenes that compare our modulated gradient scaling (MGS) and naive gradient scaling of [60] from 2-view, 4-view, and 6-view settings in TABLEXXIIIXXIV, and XXV, respectively.
D-D World Coordinate System Experiment for Entire Scenes
We present the quantitative comparison results for entire scenes that include MGS experiments on the world coordinate system from 2-view, 4-view, and 6-view settings in TABLEXXVIXXVII, and XXVIII, respectively. The results demonstrate the ability of our modulated gradient scaling (MGS) as geometric guidance.
TABLE XIII: Quantitative results for the entire scene of synthetic and real scenes from 2-view settings. Each color shading represents the best, second best and third best result, respectively.
@@ D-D World Coordinate System Experiment for Entire Scenes
TABLE XIV: Quantitative results for the entire scenes of synthetic and real scenes from 4-view settings. Each color shading represents the best, second best and third best result, respectively.
@@
TABLE XV: Quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 6-view settings. Each color shading represents the best, second best and third best result, respectively. DP-kernel and DN-kernel denote the kernel of [2] and [1].
@@
TABLE XVI: Ablation quantitative results for the entire scenes of synthetic and real scenes from 2-view settings. Each color shading represents the best, second best and third best result, respectively.
@@
TABLE XVII: Ablation quantitative results for the entire scenes of synthetic and real scenes obtained from 4-view settings. Each color shading represents the best, second best and third best result, respectively.
@@
TABLE XVIII: Ablation quantitative results for the entire scenes of synthetic and real scenes obtained from 6-view settings. Each color shading represents the best, second best and third best result, respectively.
@@
TABLE XIX: Quantitative results of the RegNeRF [6] with and without the blur kernel from 2-view to 10-view settings. For comparison, the blur kernel of the DP-NeRF [2] is employed, which is denoted as DP-kernel. Color shading represents the better result.
@@
TABLE XX: Entire experimental results for the complex optimization issue in both synthetic and real scenes from 2-view settings. Each color shading represents the best and second results for each experimental setting, respectively.
@@
TABLE XXI: Entire experimental results for the complex optimization issue in both synthetic and real scenes from 4-view settings. Each color shading represents the best and second results for each experimental setting, respectively.
@@
TABLE XXII: Entire experimental results for the complex optimization issue in both synthetic and real scenes from 6-view settings. Each color shading represents the best and second results for each experimental setting, respectively.
@@
TABLE XXIII: Quantitative comparison between naive gradient scaling of [60] and our MGS for entire scenes of synthetic and real scenes from 2-view settings. The experimental results are presented according to the two types of kernel we utilize in this paper. Color shading represents the better result.
@@
TABLE XXIV: Quantitative comparison between naive gradient scaling of [60] and our MGS for entire scenes of synthetic and real scenes from 4-view settings. The experimental results are presented according to the two types of kernel we utilize in this paper. Color shading represents the better result.
@@
TABLE XXV: Quantitative comparison between naive gradient scaling of [60] and our MGS for entire scenes of synthetic and real scenes from 6-view settings. The experimental results are presented according to the two types of kernel we utilize in this paper. Color shading represents the better result.
@@
TABLE XXVI: Quantitative comparison results for entire scenes from a 2-view setting that include MGS experiments on the world coordinate system. Color shading represents the best and second result.
DP-kernel $\times$ 15.17 .2710 .6140 13.48 .2441 .6758 15.24 .1592 .8082 15.00 .3469 .4822 13.43 .2859 .5489 14.46 .2614 .6258
DP-kernel $+$ Naive [60] 15.10 .2599 .6074 13.70 .2598 .6662 15.41 .1729 .7816 14.85 .3403 .4823 13.00 .2566 .5643 14.41 .2579 .6204
DP-kernel $+$ MGS 15.30 .2682 .6026 13.12 .2199 .6755 15.82 .1592 .6917 14.97 .3626 .4710 13.25 .2723 .5460 14.49 .2564 .5974
Real Scene Gradient Scaling Ball Basket Buick Coffee Decoration
DP-kernel $\times$ 19.25 .4315 .5415 13.59 .2255 .5975 10.86 .0918 .6892 16.06 .4164 .6220 11.06 .0899 .7539
DP-kernel $+$ Naive [60] 19.38 .4441 .5298 13.23 .2343 .5941 11.02 .1003 .6903 15.80 .3780 .6336 10.91 .0834 .7609
DP-kernel $+$ MGS 19.41 .4349 .5261 16.47 .3854 .4621 14.87 .3050 .4806 20.31 .6097 .3676 11.57 .1617 .6643
Real Scene Gradient Scaling Girl Heron Parterre Puppet Stair Average
DP-kernel $\times$ 9.41 .0744 .7648 11.73 .0565 .6825 14.10 .2013 .7586 12.72 .1821 .6405 13.87 .0276 .6817 13.26 .1797 .6732
DP-kernel $+$ Naive [60] 9.35 .0906 .7656 11.64 .0645 .6473 13.97 .1795 .7477 12.64 .1865 .6280 14.10 .0280 .6827 13.20 .1789 .6680
DP-kernel $+$ MGS 11.67 .2883 .5683 12.78 .1212 .5570 16.09 .2590 .5353 13.20 .2016 .6082 15.80 .1127 .5924 15.22 .2880 .5362
TABLE XXVII: Quantitative comparison results for entire scenes from a 4-view setting that include MGS experiments on the world coordinate system. Color shading represents the best and second result.
DP-kernel $\times$ 16.61 .3544 .5434 25.00 .7541 .1967 15.31 .1878 .8088 17.57 .4957 .3911 17.81 .4852 .4055 18.46 .4554 .4691
DP-kernel $+$ Naive [60] 16.70 .3570 .5459 25.11 .7578 .2058 15.15 .1751 .7878 17.50 .4898 .3982 17.54 .4732 .4153 18.40 .4506 .4706
DP-kernel $+$ MGS 16.71 .3612 .5380 24.48 .7441 .2072 16.63 .1934 .5907 16.90 .4590 .4038 17.96 .4887 .4105 18.54 .4493 .4300
Real Scene Gradient Scaling Ball Basket Buick Coffee Decoration
DP-kernel $\times$ 22.39 .5417 .3903 19.18 .4771 .4105 14.51 .3651 .4723 21.86 .6689 .2854 11.54 .1274 .7296
DP-kernel $+$ Naive [60] 22.84 .5624 .3842 18.70 .4611 .4302 16.33 .3849 .4132 21.05 .6320 .3154 11.17 .1222 .7454
DP-kernel $+$ MGS 21.02 .4775 .4195 19.63 .4935 .4014 16.82 .4307 .4087 23.70 .7215 .2480 16.45 .3778 .4574
Real Scene Gradient Scaling Girl Heron Parterre Puppet Stair Average
DP-kernel $\times$ 15.95 .5447 .3868 16.87 .3079 .3339 16.35 .2726 .5563 14.81 .2771 .4870 15.76 .0964 .5175 16.92 .3679 .4570
DP-kernel $+$ Naive [60] 16.53 .5400 .3763 17.30 .3133 .3287 15.72 .2485 .5773 15.30 .2899 .4753 15.60 .0998 .4976 17.05 .3654 .4544
DP-kernel $+$ MGS 16.31 .5390 .3694 17.04 .3202 .3195 16.70 .2837 .4821 15.01 .2914 .4659 16.39 .1255 .4470 17.91 .4061 .4019
TABLE XXVIII: Quantitative comparison results for entire scenes from a 6-view setting that include MGS experiments on the world coordinate system. Color shading represents the best and second result.
DP-kernel $\times$ 17.40 .3972 .4993 24.76 .7471 1878 16.10 .2033 .7592 18.35 .5266 .3951 18.60 .5370 .3628 19.04 .4822 .4408
DP-kernel $+$ Naive [60] 17.36 .3929 .5039 24.27 .7370 .2001 16.33 .2201 .7567 18.67 .5289 .3888 18.42 .5363 .3628 19.01 .4830 .4425
DP-kernel $+$ MGS 17.30 .3963 .4974 24.04 .7178 .1957 17.48 .2349 .5371 18.07 .5015 .3892 18.63 .5400 .3623 19.10 .4781 .3963
Real Scene Gradient Scaling Ball Basket Buick Coffee Decoration
DP-kernel $\times$ 23.97 .6166 .3404 20.79 .5512 .3421 17.01 .4657 .3825 25.20 .7517 .2152 13.76 .2104 .6388
DP-kernel $+$ Naive [60] 24.64 .6489 .3208 20.37 .5343 .3665 17.36 .4707 .3826 25.12 .7379 .2268 11.63 .1248 .7265
DP-kernel $+$ MGS 23.86 .6111 .3449 20.91 .5549 .3511 17.78 .4768 .3802 26.10 .7655 .2119 16.60 .3919 .4509
Real Scene Gradient Scaling Girl Heron Parterre Puppet Stair Average
DP-kernel $\times$ 17.39 .5636 .3255 19.35 .4523 .2940 18.55 .4203 .4556 19.77 .5354 .3428 17.96 .1929 .4032 19.38 .4760 .3740
DP-kernel $+$ Naive [60] 17.48 .5729 .3328 19.29 .4714 .2863 18.65 .4257 .4403 18.91 .4096 .3628 17.76 .1785 .3853 19.12 .4575 .3831
DP-kernel $+$ MGS 18.08 .5934 .3184 18.37 .4190 .3011 18.91 .3998 .3986 19.63 .5319 .3443 17.69 .1798 .3682 19.79 .4923 .3470